### v1

# Data Augmentation is Dead, Long Live Data Augmentation

Abstract
Textual data augmentation (DA) is a prolific field of study where novel techniques to create artificial data are regularly proposed, and that has demonstrated great efficiency on small data settings, at least for text classification tasks. In this paper, we challenge those results, showing that classical data augmentation is simply a way of performing better fine-tuning, and that spending more time fine-tuning before applying data augmentation negates its effect. This is a significant contribution as it answers several questions that were left open in recent years, namely : which DA technique performs best (all of them as long as they generate data close enough to the training set as to not impair training) and why did DA show positive results (facilitates training of network). We furthermore show that zero and few-shot data generation via conversational agents such as ChatGPT or LLama2 can increase performances, concluding that this form of data augmentation does still work, even if classical methods do not.
Keywords: NLP, Classification, Data Augmentation, ChatGPT, Llama2
[…]
Data Augmentation is Dead, Long Live Data Augmentation
Frédéric Piedboeuf, Philippe Langlais
Université de Montréal, RALI
frederic.piedboeuf@umontreal.ca
Abstract content
1.   Introduction
Data augmentation (DA) consists in generating artificial data points with the hope of improving the training of a model. More specifically, interpretable textual DA takes interest into whether new sentences can increase performances, with the most common task for testing algorithms being text classification. Research on the subject generally finds that DA provides a great increase on small data settings Chen et al. (2021); Kumar et al. (2021), a small increase in classification with medium datasets (up to 1000 examples) Karimi et al. (2021); Liesting et al. (2021), and almost no increase on large datasets Kobayashi (2018); Yang et al. (2020).
DA algorithms proposed in the literature have an important effect not only because they guide practitioners towards one technique or the other, but also because they contribute to the understanding of how neural networks learn. Showing, for example, that a word substitution algorithm helps a pretrained model like BERT has important implications about its inner working.
In this paper, we show that inadequate fine-tuning of the baselines (model trained without DA) from previous studies lead to an overestimation of the impact of textual DA. In fact, we show that by simply playing with the patience parameter as well as occasionally adding label smoothing Szegedy et al. (2015), we can raise the performance of the baseline classifier to be as good as performances with DA, and that when the classifier is trained correctly DA does not bring improvements. We further compare novel DA methods using ChatGPT and Llama2 and show that the only thing that can consistently increase classification performance is generating data akin to external data (with zero or few-shot data generation), rather than generating new data from the current training distribution (by paraphrasing or modifying a sentence from the training set), as classical DA approaches do. Finally, we consider more realistic approaches to DA, where we either don’t have access to clean validation data, or where we adjust the training/validation split to better reflect real-world conditions.
Overall, this paper covers several important contributions. The first and most important of them lies in showing that classical data augmentation (DA methods from before the advent of Very Large Language Models, or VLLMs) does not work on textual classification, and we also answer why previous studies showed positive results. These positive results have been a big question in recent years, as nobody could explain with satisfaction why generated sentences helped transformers learn better, the best hypothesis being that it brought some kind of regularization to the network (Feng et al., 2021; Queiroz Abonizio and Barbon Junior, 2020).
The second contribution is showing that by using VLLMs to generate novel sentences (and not just paraphrase or modify existing ones), the performance increases, but that this is still not as efficient as collecting and annotating novel data. We furthermore analyze two strategies for generating novel data, namely zero-shot data generation (using the task description) and 3-shot data generation (giving three examples to the VLLM), showing that in most case the 3-shot strategy largely outperforms zero-shot generation.
The third contribution joins a small but growing literature comparing VLLMs in showing that ChatGPT outperforms Llama2 (Liu et al., 2023; Plátek et al., 2023; Guo et al., 2023). To our knowledge, we are the first to study this in the context of data augmentation.
Finally, our last contribution is a questioning of the use of the validation data in data augmentation for small data settings. DA research often supposes the availability of clean validation samples, but its usage is not realistic in most cases. With the splits used in those studies, one finds itself with a few data points for training the model (generally less than 30), and a few hundreds for validation. In this paper, we attempt to consider better splitting and more realistic use of data.
The paper is organized as such. In Section 2, we go over the literature on interpretable textual DA. Sections 3 and 4 explain the strategies, protocol, and datasets used in this paper. Then Section 5 presents the results and Section 6 go over analysis, including the evaluation of more realistic uses of the validation data. We finally include a discussion about the meaning of our work in Section 7.
2.   Related Work
We may categorize DA techniques in three broad families : word-level operations, paraphrasing, and generative DA. Furthermore, we refer to DA techniques that were developed before VLLMs as classical DA.
The first family creates variations of existing sentences by affecting individual words. The most seminal technique of this category is EDA, which considers four operations : word substitution, word deletion, word swapping, and insertion of related vocabulary (Wei and Zou, 2019; Liesting et al., 2021). For insertion and substitution, EDA uses WordNet to find synonyms of words from the input sentences. Another simple but efficient word-level technique is AEDA Karimi et al. (2021), where random punctuation is inserted in between words of the sentence.
Other ways of replacing words have also been considered, such as the use of a pre-trained neural network to predict masked words Kobayashi (2018), using pre-trained embeddings to find words close in the embedded space Marivate and Sefara (2020), or using BERT/BART conditioned on the class to predict masked words or spans Wu et al. (2019); Kumar et al. (2021). These techniques are generally easy to implement but bring little diversity to the dataset.
The second family of techniques acts at the sentence level, by considering the whole sentence for creating paraphrases. The seminal technique representing this family of techniques is Back-translation (BT), in which a sentence is translated to another language and then back into English (Hayashi et al., 2018; Yu et al., 2018; Corbeil and Ghadivel, 2020; Edunov et al., 2018). The use of generative models for paraphrasing has also been considered, for example by encoding and decoding a sentence through a VAE (Mesbah et al., 2019; Yerukola et al., 2021; Nishizaki, 2017). Some other strategies that have been considered are modifying syntactic trees Coulombe (2018), fine-tuning BART on an in-domain corpus of paraphrases Okur et al. (2022), using an off the shelf T5 model for paraphrasing (T5-Tapaco) Piedboeuf and Langlais (2023), or using VLLMs such as ChatGPT for paraphrasing Fang et al. (2023).
Finally, generative methods aim to generate novel sentences from the same distribution as the training data. With an ideal generative strategy, this is equivalent to collecting new data, without the annotation cost associated to the collection process. Early studies of these techniques used GPT-2 (Kumar et al., 2020; Liu et al., 2020; Yang et al., 2020; Bayer et al., 2023) or VAEs Qiu et al. (2020); Piedboeuf and Langlais (2022), but recently focus has shifted to the use of VLLMs, first with GPT3 (Yoo et al., 2021; Sahu et al., 2022), and then with ChatGPT for zero or small data generation (Møller et al., 2023; Ubani et al., 2023; Shushkevich and Cardiff, 2023; Sharma and Feldman, 2023).
Some recent works pertinent to this paper are the research of Kim et al. (2022); Zhu et al. (2023)Kim et al. (2022) question the training-validation splitting methodology in semi-supervised learning, showing that it is more efficient to fine-tune on augmented samples (created with DA) and use the original training sentences as validation data, instead of using a classical training/validation split. In Zhu et al. (2023), the authors note that weakly supervised learning approaches are evaluated by assuming the availability of clean validation samples, which is not often the case when working with small data. They notably develop novel methods for testing training under more realistic small data learning settings. In Section 6, we take inspiration of those papers to define our new strategies.
3.   Data augmentation strategies
Kumar et al. (2021); Piedboeuf and Langlais (2023) compared several DA methods, from which we select strategies that have been shown to be efficient. Concretely, we test EDA, AEDA, CBERT, CBART, T5, BT, and three strategies with VLLMs : paraphrasing, zero-shot generation, and k-shot generation. All strategies are illustrated in Figure 1.
EDA (multiple operations on the words) and AEDA (insertion of punctuation) can be implemented fairly simply. Following experiments and results from the literature, we affect 10% of the words of a sentence in EDA, and use the formula given in Karimi et al. (2021) to calculate the number of punctuation signs to insert for AEDA.
CBERT and CBART both use the task of masked words prediction to generate new words conditionally on the class. Concretely, we mask words, prepend the class to the sentence (with a separation token), and fine-tune the model to predict these masked words. For generating new sentences, we follow the same process. The difference between CBERT and CBART is that CBART can predict spans of words instead of just individual tokens, allowing it more flexibility in the generated sentences.
Both BT and T5-Tapaco aim to produce paraphrases of the original sentences to bring diversity to the training set. In BT, paraphrases are generated by translating the sentence into a second language and then back into English, and we use WMT [^1] https://huggingface.co/facebook/wmt19-De-en with German as a pivot language, which has shown good performances in the past Edunov et al. (2018). T5-Tapaco makes use of T5-small-Tapaco [^2] https://huggingface.co/hetpandya/t5-small-tapaco, which is a T5 model fine-tuned on the TaPaCo paraphrase corpus (Scherrer, 2020), allowing us to directly generate paraphrases of the sentences.
In terms of VLLMs, we test three strategies, which are the three standards strategies we could find in the literature : paraphrasing existing sentences, zero-shot generation by giving the VLLMs a description of the task, and a k-shot generation strategy where we give examples of the given class. We give the exact prompts used in appendix A. Furthermore, we test the use of Llama2 (using LLama-2-13B-Chat) with the same prompts. We denote in this paper these strategies by the letters P (paraphrase), ZS (zero-shot), and 3S (3-shot generation). We opted on 3-shot for the K-shot generation strategy based on results from exploratory experiments.
Figure 1: Strategies tested in this paper.
Finally, we implement three simple baselines. The first one consists in training the network without the use of DA (denoted “Baseline”), and the second one is an idealized strategy (denoted “‘Perfect”) where we fetch additional unused data from the training set to act as generated data. This gives an idea of the results obtainable should one collect and annotate data instead of generating new data.
The last baseline is a strategy which we denote “Copy”, and which artificially inflates the size of the training set by copying multiple times the original data, without modification. If the only effect of DA is truly to help fine-tune the network better, and the modification brought by the DA strategies are not helpful for that, then the copy strategy should be just as efficient as all other classical data augmentation strategies. We give the code with all hyperparameters in the associated github. [^3] Github will be made available upon acceptance of the paper.
4.   Datasets and protocol
We use five datasets to test the various data augmentation strategies : SST-2, Irony, IronyB, TREC6, and SNIPS. SST-2 (Socher et al., 2013) is a dataset of movie review classification with the binary classes of positive or negative. Irony and IronyB (Van Hee et al., 2018) are the binary and multiclass version of an Irony detection dataset. In the binary task (Irony), one must detect if tweets are ironic or not, and in the multiclass version one must as well determine which type of Irony the tweet represents, if the tweet is ironic (between polarity clash, situational irony, and other irony). TREC6 (Li and Roth, 2002) is a task of question classification, where questions must be separated into six classes (abbreviation, description, entities, human beings, locations, and numeric values). Finally, SNIPS Coucke et al. (2018) is a dataset of intent classification, where short commands have to be classified into different intent (PlayMusic, GetWeather, RateBook, SearchScreeningEvent, SearchCreativeWork, AddToPlaylist, BookRestaurant). Some characteristics of the datasets are available in Table 1.
Table 1: Characteristics of the tasks tackled in this study. The length of the sentences is defined by the number of white-space separated tokens.
Name SST2 Irony IronyB TREC6 SNIPS
|classes| 2 2 4 6 7
|train| 6920 2683 2681 5452 13084
len. sents. 19.3 13.7 13.7 10.2 9.3
|val| 872 460 460 500 700
|test| 1821 3834 3832 492 700
To define our training protocol, we first look at the literature to understand what is usually done. Piedboeuf and Langlais (2023) use variable patience for fine-tuning, depending on the dataset and dataset size, between 5 and 20 epochs, and Kumar et al. (2021) use 100 epochs of warmup followed by 8 epochs of fine-tuning from which they select the best model. Wei and Zou (2019) use early stopping with patience of 3 epochs on CNN and RNNs.
In contrast, we use a patience parameter of 50, using the validation set to find when to stop, and we fine-tune the learning rate and the label smoothing (a mean of introducing noise to the labels for regularization) before applying data augmentation, using grid search. We define here the patience as the number of epochs during which the validation performance doesn’t increase before we stop the training process, and we then select the best model from previous epochs. We report accuracy for binary tasks and the macro-F1 for multiclass tasks.
Our focus in this paper is data augmentation for small data learning, for which we use starting size of 10 and 20, but we also report results for medium training sizes (500 and 1000) in Appendix B, a setting which is commonly explored in the literature. When subsampling to create the training set, we make sure to choose an equal number of data points for each class, by going over the fixed dataset size if needed (for example, the actual dataset size for TREC6 and starting size of 10 is 12; two sentences per class).
The final parameter is the ratio of generated-to-genuine examples. There seems to be no strict rules about this ratio, but we found in both the literature and experiments that smaller sizes of datasets benefited more from larger ratios, while larger starting sizes did not show improvements with ratios larger than one. Therefore, we use a ratio of 10 for the small data settings and of one for the larger dataset sizes. [^4] Based on our main results, we can hypothesize that this is simply due to pretrained models needing more time to be fine-tuned correctly on smaller datasets than medium or large ones, and the larger ratio leads to a larger number of batches, meaning more training time.
Something worth noting is that all of these protocols (including ours) use the validation data to fine-tune both the DA algorithm and the model to train. This is somewhat questionable, as keeping only 10 or 20 data points out of a few hundreds might not be the best way to utilize the data. While we keep this protocol in the main section to stay in accordance with the DA literature, we propose more realistic evaluation protocol of DA in Section 6.
The norm in data augmentation literature is to give the average and standard deviation over a fixed number of runs. However, to confirm whether the occasional improvements are due to randomness of the model or the effect of DA, we include in this paper statistical tests.
Table 3: Average metric over 15 runs for the training set sizes of 10 (left) and 20 (right) with a ratio of 1. We report accuracy for binary tasks and macro-f1 for multiclass ones. STDs are between 0.6 and 3.0, depending on the dataset. Stars represent results for which the difference with the baseline was found to be statistically significant.
SST2 Irony IronyB Trec6 SNIPS Average
Baseline 62.4/67.5 54.1/61.9 25.8/29.8 34.4/43.3 78.5/80.4 51.1/56.6
Perfect 82.3*/85.8* 62.4*/58.7 19.3/33.7 72.5*/80.6* 94.0*/94.9* 66.1/70.7
Copy 61.0/68.1 54.4/57.1 25.5/30.3 36.4/49.3* 77.7/83.4* 51.0/57.6
EDA 63.4/69.0 54.9/58.9 23.8/30.1 35.8/48.0* 76.6/83.2* 50.9/57.8
AEDA 62.4/67.1 55.4/56.9 23.4/27.4 33.2/45.7 78.0/84.0* 50.5/56.2
BT 60.2/66.8 55.1/58.0 24.7/27.9 29.1/51.5* 75.0/82.6* 48.8/57.4
CBERT 62.5/65.4 56.3*/58.5 23.2/28.3 36.4/44.5 76.3/83.4* 50.9/56.0
CBART 62.5/65.6 54.5/57.7 23.6/29.0 36.9/47.2 77.0/84.5* 50.9/56.8
T5 64.1/67.4 56.3/58.3 24.1/28.1 34.3/48.5* 76.7/84.1* 51.1/57.2
GPT3.5-P 65.0*/68.3 56.6*/57.5 22.7/28.3 34.0/40.5 80.2/84.7* 51.7/55.9
GPT3.5-ZS 82.0*/78.7* 56.4*/56.1 23.5/24.6 40.0*/48.1* 84.8*/88.7* 57.3/59.2
GPT3.5-3S 87.7*/76.1* 49.2/53.6 20.9/24.6 58.4*/63.9* 87.2*/89.6* 60.7/61.5
Llama2-P 66.3*/68.6 55.3/57.1 24.6/24.6 33.4/41.6 78.1/82.8* 51.6/55.0
Llama2-ZS 74.8*/74.0* 54.3/57.2 22.2/24.1 37.3/47.3 78.1/84.1* 53.4/57.3
Llama2-3S 63.3/64.4 54.1/54.8 21.0/27.2 48.6*/55.8* 83.3*/88.2* 54.0/58.1
For each starting size (10, 20) and dataset, we perform a two-tailed paired t-test, comparing the algorithms to each other [^5] We choose to do this instead of a Hotelling’s T-square to take into account the fact the DA might work on some datasets but not others. For example, one would intuitively expect synonym replacement to work better on sentiment classification than on Irony detection.. As is standard, we assess the two distributions to be different if the p-value is lower than 0.05.
5.   Results
We first show the inefficiency of the fine-tuning protocol in past DA papers by reporting the results they obtain without data augmentation and comparing it to our results, without data augmentation as well. We attempt to replicate the starting sizes, changing only the fine-tuning protocol to see the difference a longer patience, label smoothing, and better fine-tuning does.
Reported Ours
Kumar + Ubani SST-2 52.9 60.6
Kumar + Ubani SNIPS 48.6 66.8
Kumar + Uabni TREC6 79.4 91.6
Piedboeuf SST-2 87.7 87.7
Piedboeuf Irony 65.6 67.0
Piedboeuf IronyB 42.4 43.9
Piedboeuf TREC6 81.0 84.5
Table 2: Fine-tuning comparison from the literature and in this paper. All results are from training BERT without data augmentation, showing the difference made by using a longer patience as well as label smoothing.
Table 2 presents the results of our experiments compared to results from the literature. As we can see, having a larger patience parameter and applying label smoothing increase the performance significantly, gaining between 2 and 10%. This is the core of our argument, as our hypothesis is that the gains observed in previous studies stems only from data augmentation, giving more time for the model to learn correctly. Furthermore, if the DA strategies involve simple transformations of the sentences, it is likely that BERT is not learning new information from these transformations, and that the same result will be achieved with the Copy strategy.
We now turn to DA algorithms, showing that using the updated training protocol does not result in higher gain, and that indeed the supposed gains from DA that were shown in the literature can more likely be explained by inadequate training of the classifier. Results are presented in Table 3.
For the small data setting, we notice that, on average, only two strategies really outperforms the baseline : zero-shot and 3-shot generation, both using VLLMs. Of these two strategies, the 3-shot generation outperforms the zero-shot one, but not on all datasets. We explore this phenomenon in Section 6.2.
An important thing to note is that while some algorithms perform statistically better than the baseline at times, so does the “Copy” strategy, which is a simple duplication of the data. As the duplication of data should bring no new information, this suggests that the training protocol was suboptimal and that the difference between the baseline and the DA results is due to this rather than a real effect of the DA algorithm. One avenue that could be explored to bridge the gap would be to use sampling with replacement while fine-tuning BERT, which should be equivalent to duplication of data [^6] There would still be minor differences as sampling with replacement would not represent each sample equally, but this difference should reasonably not have a large impact.. The results for medium sizes (500 and 1000) are presented in Appendix B, and are concordant with more recent literature showing a lack of improvement on such a setting.
We give in Figure 4 the results of the statistical tests. To simplify reading, we regroup together all tests between the same algorithms, meaning all tests between algorithm A and algorithm B are compiled together. Each entry in the heatmap is therefore composed of 10 t-tests (5 datasets * 2 starting sizes), and we report the percentage of t-tests for which distributions were found to be statistically different, counting only the entries for which the row algorithm performs better than the column algorithm. This gives, for each row, the direct measure of how many times it beats the algorithm in the column by a statistically significant margin. [^7] We provide in the code the full saved results for all experiments, as well as the code necessary to run all statistical tests.
Baseline
Perfect
Copy
EDA
AEDA
BT
CBERT
CBART
T5
GPT3.5-P
GPT3.5-ZS
GPT3.5-3S
Llama2-P
Llama2-ZS
Llama2-3S
Baseline
0
20
10
10
10
30
10
10
10
20
20
40
10
30
30
Perfect
70
0
70
70
70
70
70
70
70
70
70
80
80
80
80
Copy
20
10
0
0
10
20
10
0
0
30
10
40
20
20
30
EDA
20
10
20
0
10
20
0
0
0
10
20
40
20
10
20
AEDA
10
10
0
0
0
20
0
0
0
10
0
30
20
0
0
BT
20
10
0
0
10
0
10
10
0
10
10
30
10
10
20
CBERT
20
10
10
0
10
10
0
10
0
10
0
20
10
0
20
CBART
10
10
0
0
0
10
0
0
0
10
0
30
10
10
10
T5
20
10
0
0
0
20
10
0
0
10
0
40
10
10
20
GPT3.5-P
30
10
10
20
10
40
20
20
10
0
10
30
10
20
20
GPT3.5-ZS
70
10
60
50
50
50
60
50
50
60
0
40
60
50
30
GPT3.5-3S
60
10
60
60
60
60
60
60
60
60
40
0
60
60
50
Llama2-P
20
10
10
0
10
20
20
10
0
0
0
30
0
0
10
Llama2-ZS
30
10
20
20
20
30
20
20
20
30
0
20
40
0
30
Llama2-3S
40
0
40
40
40
40
40
40
40
40
20
20
40
40
0
Figure 2: Percent of times the row algorithm performs statistically better than the column algorithm, with a p-value threshold of 0.05 and using a two-tails paired t-test, and across the two small data settings (10/20).
From the statistical tests, we note that only three methods outperform the others by a significant and consistent margin : the perfect strategy, in which we simulate collecting and annotating external data, and the zero/k-shot generation with ChatGPT. All others perform equivalently and do not beat the baseline most of the time, or at least not by a significant margin. This suggests that the slight random variations we observe, which could be taken as a sign that some algorithm performs better than other, are due to randomness of the network and not a superior performance.
6.   Analysis
The preceding section evaluated data augmentation algorithms using a standard training protocol. While our main result is that the efficiency of DA is generally over-estimated, this does not imply data augmentation is never useful. In this section, we aim to test more realistic settings with the validation data, inspired by Kim et al. (2022); Zhu et al. (2023), who study the same problems but in different settings (semi-supervised learning and weakly supervised learning) and then discuss what our results mean for textual DA. While Kim et al. (2022); Zhu et al. (2023) findings are relevant to ours, our interest in the results differ. In both papers, the authors attempt to find the best way to use the available data. In our case, we test several settings we think are more realistic, but the point of interest is to find if data augmentation is useful when there is little or no validation data for fine-tuning the classifier. Indeed, there is little doubt that using validation data as training data will be more efficient than the current accepted setting.
6.1.   Evaluating small data learning with more realistic use of data
In academic papers, DA is often evaluated with the assumption that we have access to clean validation data, often in larger quantity than the training data. This validation data is used to fine-tune both the DA algorithm and the classifier, which results in unrealistic settings for practitioners, who may not have access to validation data or want to use their data in better ways. In this section, we evaluate the use of DA in various low-data settings in which we redefine the training and validation split, using different strategies.
We design experiments using two setups, electing to test the setups with the starting size of 10. This is motivated by the long time it takes to fine-tune and run all algorithms [^8] Depending on the final training set size and the DA algorithms, running one experiment takes between 30 and 75 minutes, excluding the grid-search for the hyperparameters, on a RTX 3090., and the fact that the efficiency of DA was the same for both starting sizes, making duplication of the experiments unnecessary. An open question is how these settings would work if we had more initial training data, but we leave this for future work.
Firstly, we assume the validation data is not available (noVal), and we only have training data. In this setting, we train for a random number of epochs between 50 and 150, which is the range for which our models performed best in small data learning.
Secondly, we assume the presence of validation data but redefine the training/validation split, by either 1- using it all as training data and training for a random number of epochs between 4 and 8 (valAsTrain), or 2- keeping some of it as validation data (10, 250) to fine-tune both DA algorithms and the classifier (10 val/250 val).
Overall, this gives us four settings to test, illustrated in Figure 3. We select the following algorithms : Copy, Perfect, EDA, AEDA, BT, CBERT, BART, T5, and GPT3.5-ZS. We use a limited selection for efficiency and because the goal of these experiments is simply to establish whether DA helps at all, rather than which one works best.
Figure 3: Graphic representation of the four settings we test for data augmentation on small data learning. Blue represent the original training set, purple the validation set, and yellow, the test set.
Finally, to keep in line with other data augmentation studies, we use a ratio of 10 when the final training size is small (noVal) and of 1 for the other three settings.
noVal valAsTrain 10 val 250 val
Baseline 47.9 75.8 64.6 64.7
Perfect 68.8 82.4 69.4 69.9
Copy 49.6 78.6 66.3 65.8
EDA 49.1 77.8 66.0 64.9
AEDA 49.4 77.6 66.0 66.3
BT 44.8 75.9 64.6 65.0
CBERT 49.7 76.8 65.3 63.9
BART 50.6 76.6 65.3 63.9
T5Par 49.4 77.2 65.3 64.6
GPT3.5-ZS 55.9 76.9 65.6 64.4
GPT3.5-3S 57.2 71.6 67.2 64.8
Table 3: Average metrics over the five datasets for different settings and DA strategies. The best result for each setting is in bold.
Results are presented in table 3, and several important observations can be made. First, in almost all cases and contrary to the main results of this paper, data augmentation reveals itself to be useful. However, the fact that the “Copy” strategy often outperforms the other reinforces our belief that the use of data augmentation is simply to facilitate fine-tuning the network.
Second, there is no significant difference in performance between having 10 and 250 data points as validation data, with the rest as training set. This could be related to diminished increase in model performance as dataset size augment, or point that there is a trade-off in validation/training size and that a validation set of 10 data points is too small to be of use.
Third, we confirm here that methods adding something akin to external data (GPT3.5-ZS and GPT3.5-2S) are the most useful techniques, but only on small training sizes. On larger training sizes, they don’t bring any improvement whatsoever to the model, most likely due to the general inefficiency of data augmentation in this setting. [^9] This is concordant with the literature and our own experiments. We also note that while the performance of GPT3.5-ZS is higher than GPT3.5-3S for two settings, this is due exclusively to its poor performance on the Irony datasets, which we discuss in Section 6.2. For the other three datasets, GPT3.5-3S outperforms GPT3.5-ZS by a large margin.
Finally, it seems that the best strategy overall when the total amount of available data is only a few hundred is to use all of them as training data, with the Copy strategy. This outperforms by a large margin the strategies where we keep some data as validation data, most likely because these strategies overfit. As mentioned, however, there is a large gap between 10 and 250 data and it is possible that some amount of data in the middle would perform best. We leave this for future work.
6.2.   Analysis of the VLLMs
Our main results showed that in a small data setting, 3-shot generation was often more efficient than zero-shot generation, by as much as 18% in some cases. However, this does not hold for the two irony datasets, for which zero-shot outperforms 3-shot. In this section, we analyze the output of the algorithms, in the hope of understanding better why some prompt mechanisms work well on some datasets but not on others.
We first show examples of generated sentences in Tables 4 and 5. For the sake of space, we only show generated sentences for the classes of SST-2 and Irony, but we give all sentences generated by ChatGPT in the associated Github.
Negative
Dataset apparently reassembled from the cutting-room floor of any given daytime soap.
GPT3.5-P It appears to have been pieced together from discarded scenes of any typical daytime soap opera.
GPT3.5-ZS The film was a major disappointment, lacking any coherent plot or engaging characters.
GPT3.5-3S It’s a shame that such a promising concept was wasted on a poorly executed film.
Positive
Dataset béart and berling are both superb , while huppert … is magnificent.
GPT3.5 - P Both Béart and Berling deliver exceptional performances, while Huppert… she is truly magnificent.
GPT3.5-ZS This film is a refreshing breath of fresh air, offering a unique and original take on familiar themes.
GPT3.5-3S This film is a delightful and heartwarming celebration of friendship, full of laughter and touching moments.
Table 4: Examples of generated sentences for the SST-2 dataset.
Non-Ironic
Dataset Couldn’t have asked for a better night.
GPT3.5-P Could not have hoped for a more perfect evening.
GPT3.5-ZS Wow, I absolutely love it when my phone battery dies in the middle of an important call. Just what I needed!
GPT3.5-3S Enjoying a relaxing day at the beach with clear skies and warm sunshine. #PerfectDay
Ironic
Dataset work should be fun today :unamused_face:
GPT3.5 - P Today at work is expected to be unenjoyable :unamused_face:
GPT3.5-ZS Oh, how wonderful! Another day of unexpected surprises and setbacks. It’s like the universe’s way of keeping me on my toes!
GPT3.5-3S Declared I’d be more organized and lost my keys for the umpteenth time. I’m the queen of tidiness! :key::mag: #OrganizationQueen
Table 5: Examples of generated sentences for the Irony dataset.
By observing generated sentences, we see that ChatGPT has difficulty understanding the concept of Irony, as well as sticking to the more familiar language of Twitter. As such, adding more sentences doesn’t help because the new sentences are too far from the training distribution to bring valuable information. Nevertheless, the 3-shot strategy does seem to bring the generated sentences closer to the training distribution, as we can at least observe that the new sentences contain hashtags and emojis.
7.   Discussion
In this paper, we test data augmentation for sentence classification and show that in both medium and small data learning, performances of DA had been overestimated by the use of inadequate training protocols. Furthermore, while we have shown that the classical methods of DA are inefficient, we have also demonstrated that simulating external data collection with VLLMs does improve the performance, and future work should therefore focus on this rather than techniques modifying the sentences.
We want to emphasize that our results are only looking at balanced classification task, and furthermore only at short text classification using BERT as a classifier. While this seems a small field of study, it is one of the most popular in textual DA literature, making our findings significant.
Future work should also evaluate whether the protocols are adapted for fields close to interpretable DA. Notably non-interpretable DA, or DA for unbalanced data, might still be an efficient mean of increasing performances. Other textual tasks might also still benefit from DA, due to different goals. As an example, data augmentation for the task of keyphrase generation (Ray Chowdhury et al., 2022; Garg et al., 2023) aims to encourage the network to generate more keyphrases that are absent from the input. It is more than plausible that data augmentation is useful for this field. As another example, data augmentation is sometime used to improve explainability (Chen and Ji, 2020; Ansari et al., 2023), and protocol should therefore be checked thoroughly in independent fields to ensure that it works as intended.
8.   Conclusion
In this paper, we look at the protocol surrounding data augmentation, and demonstrate that data augmentation is mainly used for easier fine-tuning, an effect that can be replicated by simply giving more care to the fine-tuning of the network, or by using a simple strategy where we duplicate the data. We also consider other, more realistic applications of DA, where validation data is either not available or used as training data, instead of keeping hundreds of examples for validation and only a dozen for training the model.
These findings are important as they answer many questions in textual DA which have been standing for years, namely which data augmentation algorithms is best (they all perform similarly except for data generation), why does it help (gives more time to the network to learn), and what makes a generated sentence informative (they don’t bring contradictory information to the network). As noted, data generation remains one of the best way to perform data augmentation, if one can get the VLLMs to generate data adequate for the training distribution.
Future work on textual DA should explore ways to bridge this gap and generate data which is more diverse as well as correspond more closely to the real task. Another pertinent exploration that remains to be done is to thoroughly study the difference between Llama2 and the GPT family. In this paper, we used Llama2-13B-Chat and concluded that it was less efficient than the use of ChatGPT, but it remains to be seen if this is true for all sizes of the model. We also used default parameters of ChatGPT, to stay in line with Piedboeuf and Langlais (2023). It is very probable however that other configurations are more efficient, and this should be studied as well. Finally, we attempted to evaluate DA in more realistic settings, by redefining training/validation split, but the testing should be taken further. For example, it is very likely that a jackknife approach would be a better evaluation method for DA in small data setting, and this should be considered in future research on data augmentation.
9.   Limitations
As stated on the paper, this study as a limited, although important, scope of study. We didn’t look at the impact of data augmentation on other languages than English, or other textual task, and while we performed fine-tuning to the best of our knowledge, it is possible that some combinations of hyperparameters that were not explored change the conclusions here. We also omitted an ablation study due to time limitation, but this would provide more insight into the mechanism of training BERT. Other limitations about the models are the limited hyperparameters explored for ChatGPT, the limited models used for Llama2, and the fact that we did not check the results with other classifiers than BERT.
10.   Bibliographical References
[…]
Appendix A Prompts
In this section, we describe the prompts we used for ChatGPT and LLama2. For the paraphrasing strategy, we copy Piedboeuf and Langlais (2023), who simply asks for paraphrases in batch if the ratio is 1 and for multiple paraphrases if the ratio is larger than one. We refer to their paper for more detail on the prompting.
For the zero and three-shot generation, we use the template shown in Table 6.
DATASET_DESC CLASS_DESC
movie reviews negative or somewhat negative
positive or somewhat positive
headline Fake/Real news classification Real
Fake
Ironic tweet detection Non-Ironic Tweets
Ironic Tweets
Ironic tweet detection Tweets that are not ironic
Tweets ironic by polarity contrast, where the polarity is inverted between the literal and intended evaluation
Tweets ironic by Situational Irony, where a situation fails to meet some expectation
Tweets ironic by Other type of Irony, where the Irony is neither by Polarity Contrast or by Situational Irony
Question classification Questions about an abbreviation
Questions about an entity (event, animal, language, etc)
Question concerning a description (of something, a definition, a reason, etc)
Questions about a human (description of someone, an individual, etc)
Questions about a location
Questions about something numerical (weight, price, any other number)
Table 6: Prompt patterns for the zero-shot strategies for ChatGPT and Llama. The prompt is of the form “Here are some examples of \$CLASS_DESC from a dataset of \$DATASET_DESC: \$EXAMPLES. Can you create 10 more sentences of that class that would fit the dataset ”.
We referred to the description given in the original papers of each dataset to craft informative prompts.
Appendix B Supplementary Results
Table 8: Average metric over 15 runs for the training set sizes of 500 (left) and 1000 (right) with a ratio of 1. We report accuracy for binary tasks and macro-f1 for multiclass ones. STDs are between 0.6 and 3.0, depending on the dataset. Stars represent results for which the difference with the baseline was found to be statistically significant.
SST2 Irony IronyB TREC6 SNIPS Average
Baseline 87.7/88.8 67.0/68.1 43.9/45.9 84.5/87.9 95.6/96.2 75.8/77.4
Perfect 88.9*/89.7* 67.3/71.1* 46.0*/48.5* 87.6*/89.3 96.1*/96.9* 77.2/79.1
Copy 87.7/88.8 66.2/69.2 43.7/46.7 83.8/86.3 95.5/96.4 75.4/77.5
EDA 88.1/88.9 65.9/69.4 43.6/47.2* 82.5/86.1 95.8/96.3 75.2/77.6
AEDA 88.0/89.0 65.1/69.5 42.8/46.7 82.7/87.4 96.1*/96.3 74.9/77.8
BT 88.0/89.2* 64.9/68.4 44.2/45.8 83.1/87.1 95.5/96.3 75.1/77.3
CBERT 87.6/88.6 64.8/69.5 43.6/46.1 82.9/84.7 95.4/96.3 74.9/77.0
CBART 87.7/88.7 66.0/68.8 43.6/45.1 81.2/82.5 95.9/96.2 74.9/76.3
T5 87.5/88.8 64.9/67.7 43.6/45.6 82.9/84.8 95.8/96.3 75.0/76.6
GPT3.5-P 88.2*/89.3* 65.1/69.5* 44.3/47.5 84.5/86.8 95.5/96.1 75.5/77.8
GPT3.5-ZS 87.7/88.8 64.2/67.4 42.0/45.1 81.9/86.0 95.5/95.8 74.3/76.6
GPT3.5-3S 88.1*/88.6 64.9/69.0 40.4/44.4 83.0/86.8 95.9/96.2 74.5/77.0
Llama2-P 87.8/88.9 65.4/68.1 44.1/46.1 83.3/85.9 95.7/96.4 75.3/77.1
Llama2-ZS 88.1*/88.8 65.0/69.2 42.5/45.7 82.2/86.0 95.6/96.0 74.7/77.2
Llama2-3S 87.8/88.4 65.0/67.9 42.8/46.2 83.8/85.3 95.7/96.4* 75.0/76.8
Baseline
Perfect
Copy
EDA
AEDA
BT
CBERT
CBART
T5
GPT3.5-P
GPT3.5-ZS
GPT3.5-3S
Llama2-P
Llama2-ZS
Llama2-3S
Baseline
0
0
0
20
10
10
30
20
20
0
40
30
20
30
30
Perfect
80
0
90
70
70
90
100
80
90
80
100
90
90
100
90
Copy
0
0
0
0
0
0
0
30
10
0
40
20
0
10
20
EDA
10
0
0
0
0
0
0
20
20
0
50
20
10
10
10
AEDA
10
0
0
10
0
10
20
20
20
20
40
30
20
20
30
BT
10
0
0
0
0
0
20
20
10
0
20
20
0
20
10
CBERT
0
0
0
0
0
0
0
10
10
0
30
20
10
0
10
CBART
0
0
0
0
0
0
10
0
0
0
20
10
0
0
0
T5
0
0
0
0
0
0
0
10
0
0
20
10
0
10
0
GPT3.5-P
30
0
20
20
10
10
30
40
50
0
60
30
0
40
20
GPT3.5-ZS
0
0
0
0
0
0
0
10
0
0
0
10
0
0
0
GPT3.5-3S
10
0
10
0
0
0
10
10
10
10
30
0
0
0
0
Llama2-P
0
0
0
0
0
0
10
10
0
0
30
20
0
10
10
Llama2-ZS
10
0
10
0
0
0
0
10
10
0
20
10
10
0
20
Llama2-3S
10
0
0
0
0
0
10
10
0
10
20
20
0
10
0
Figure 4: Percentage of times the row algorithm performs statistically better than the column algorithm, with a p-value threshold of 0.05 and using a two-tails paired t-test and with the starting sizes of 500/1000.