### v2

# On Evaluation Protocols for Data Augmentation in a Limited Data Scenario

Abstract
Textual data augmentation (DA) is a prolific field of study where novel techniques to create artificial data are regularly proposed, and that has demonstrated great efficiency on small data settings, at least for text classification tasks. In this paper, we challenge those results, showing that classical data augmentation (which modify sentences) is simply a way of performing better fine-tuning, and that spending more time doing so before applying data augmentation negates its effect. This is a significant contribution as it answers several questions that were left open in recent years, namely : which DA technique performs best (all of them as long as they generate data close enough to the training set, as to not impair training) and why did DA show positive results (facilitates training of network). We further show that zero- and few-shot DA via conversational agents such as ChatGPT or LLama2 can increase performances, confirming that this form of data augmentation is preferable to classical methods.
On Evaluation Protocols for Data Augmentation
in a Limited Data Scenario
Frédéric Piedboeuf Université de Montréal, RALI frederic.piedboeuf@umontreal.ca                        Philippe Langlais Université de Montréal, RALI
1 Introduction
Data augmentation (DA) consists in generating artificial data points with the hope of improving the training of a model. In this paper, we focus on interpretable textual DA (methods that generate new sentences) for text classification in a limited data scenario, a practical setting of interest Chen et al. (2021). Research generally finds that DA provides a great increase on small data settings Chen et al. (2021); Kumar et al. (2021), a small increase in classification with medium datasets (up to 1000 examples) Karimi et al. (2021); Liesting et al. (2021), and almost no increase on large datasets Kobayashi (2018); Yang et al. (2020); Okimura et al. (2022).
In this paper, we show that existing experimental protocols in DA studies are misleading. In particular, inadequate fine-tuning of the baseline (model trained without DA) from previous studies lead to an overestimation of the impact of textual DA, and training the models for longer results in an absence of gains with previous DA methods. We also consider more realistic protocols for DA evaluation, where we either don’t have access to clean validation data, or where we adjust the training/validation split to better reflect real-world conditions. Lastly, we compare newer DA methods using ChatGPT Ouyang et al. (2022) and Llama2 Touvron et al. (2023) and show that the only thing that can consistently increase classification performance is generating data akin to external data (with zero or few-shot data generation), rather than generating new data from the current training distribution (by paraphrasing or modifying training sentences), as classical DA approaches do.
Overall, this paper covers several important contributions. The first and most important of them lies in showing that classical data augmentation (DA methods from before the advent of very Large Language Models, or LLMs) does not work on textual classification, and we also answer why previous studies showed positive results. These positive results have been a big question in recent years, as nobody could explain with satisfaction why generated sentences helped transformers learn better, the best hypothesis being that it brought some kind of regularization to the network (Feng et al., 2021; Queiroz Abonizio and Barbon Junior, 2020).
Our second contribution is a questioning of the use of the validation data in data augmentation for small data settings. DA research generally supposes the availability of clean validation samples, an often unrealistic setting. With the splits used in those studies, one finds themself with a few tens of data points for training the model and a few hundreds for validation. In this paper, we consider better splitting and more realistic uses of data, showing that it is often more advantageous to use all available data as training data and to fine-tune for a longer time.
The third contribution is showing that by using LLMs to generate novel sentences (and not just paraphrasing existing ones), the performance increases, but that this is still not as efficient as manually collecting and annotating novel data. We furthermore analyze two strategies for generating novel data, namely zero-shot data generation (using the task description) and 3-shot data generation (giving three examples to the LLM), showing that in most cases the 3-shot strategy largely outperforms zero-shot generation.
The last contribution joins a small but growing literature comparing LLMs in showing that ChatGPT outperforms Llama2 (Liu et al., 2023; Plátek et al., 2023; Guo et al., 2023). To our knowledge, we are the first to study this in the context of DA.
The paper is organized as such. Section 2 goes over the literature on interpretable textual DA. Sections 3 and 4 explain respectively the datasets and the DA methods we compare. Section 5 describes our evaluation protocols, then Section 6 presents our results. We discuss our work in Section 7.
2 Related Work
We may categorize DA techniques in three broad families : word-level operations, paraphrasing, and generative DA. We note that this is not an absolute categorization, and each of these families regroup several families of techniques. We also refer to techniques that were developed before LLMs as classical DA. These modify the starting sentences in some ways, bringing variations to the dataset.
The first family does so by affecting individual words. The most seminal technique of this category is EDA, which considers four operations : word substitution, word deletion, word swapping, and insertion of related vocabulary (Wei and Zou, 2019; Liesting et al., 2021). For insertion and substitution, EDA uses WordNet Miller (1998) to find synonyms of words from the input sentence. Another simple word-level technique is AEDA Karimi et al. (2021), where random punctuation is inserted in between words of the sentence.
Other ways of replacing words have also been considered, such as the use of a pre-trained neural network to predict masked words Kobayashi (2018), using pre-trained embeddings to find words close in the embedded space Marivate and Sefara (2020), or using BERT/BART conditioned on the class to predict masked words or spans Wu et al. (2019); Kumar et al. (2021). While easy to implement, those techniques tend to bring little diversity.
The second family of techniques acts at the sentence level, by considering the whole sentence for creating paraphrases. The seminal technique representing this family is Back-translation (BT), in which a sentence is translated to another language and then back into English (Hayashi et al., 2018; Yu et al., 2018; Corbeil and Ghadivel, 2020; Edunov et al., 2018). The use of generative models for paraphrasing has also been considered, for example by encoding and decoding a sentence through a VAE (Mesbah et al., 2019; Yerukola et al., 2021; Nishizaki, 2017). Some other strategies that have been considered are modifying syntactic trees Coulombe (2018), fine-tuning BART on an in-domain corpus of paraphrases Okur et al. (2022), using an off the shelf T5 model for paraphrasing (T5-Tapaco) Piedboeuf and Langlais (2023), or using LLMs such as ChatGPT for paraphrasing Fang et al. (2023).
Finally, generative methods aim to generate novel sentences from the same distribution as the training data. With an ideal generative strategy, this is equivalent to collecting new data, without the annotation cost associated to the collection process. Early studies of these techniques used GPT-2 (Kumar et al., 2020; Liu et al., 2020; Yang et al., 2020; Bayer et al., 2023) or VAEs Qiu et al. (2020); Piedboeuf and Langlais (2022), but recently focus has shifted to the use of LLMs, first with GPT3 (Yoo et al., 2021; Sahu et al., 2022), and then with ChatGPT for zero or small data generation (Møller et al., 2023; Ubani et al., 2023; Shushkevich and Cardiff, 2023; Sharma and Feldman, 2023).
Some recent works pertinent to this paper are the research of Kim et al. (2022); Zhu et al. (2023)Kim et al. (2022) question the training-validation splitting methodology in semi-supervised learning, showing that it is more efficient to fine-tune on augmented samples (created with DA) and use the original training sentences as validation data, instead of using a classical training/validation split. In Zhu et al. (2023), the authors note that weakly supervised learning approaches are evaluated by assuming the availability of clean validation samples, which is not often the case when working with small data. They notably develop novel methods for testing training under more realistic small data learning settings. In Section 5, we take inspiration of those papers to define our new protocols.
3 Datasets
We use five popular datasets to test the various data augmentation strategies : SST-2, Irony, IronyB, TREC6, and SNIPS. SST-2 (Socher et al., 2013) is a dataset of movie review classification with the binary classes of positive or negative. Irony and IronyB (Van Hee et al., 2018) are the binary and multiclass version of an Irony detection dataset. In the binary task (Irony), one must detect if tweets are ironic or not, and in the multiclass version one must as well determine which type of Irony the tweet represents, if the tweet is ironic (between polarity clash, situational irony, and other irony). TREC6 (Li and Roth, 2002) is a task of question classification, where questions must be separated into six classes (abbreviation, description, entities, human beings, locations, and numeric values). Finally, SNIPS Coucke et al. (2018) is a dataset of intent classification, where short commands have to be classified into different intents such as PlayMusic or GetWeather. Some characteristics of the datasets are available in Table 1.
Name SST2 Irony IronyB TREC6 SNIPS
|classes| 2 2 4 6 7
sent. len. 19.3 13.7 13.7 10.2 9.3
|train| 6920 2683 2681 5452 13084
|val| 872 460 460 500 700
|test| 1821 3834 3832 492 700
Table 1: Characteristics of the classification tasks tackled in this study. The length of the sentences is defined by the number of white-space separated tokens.
Due to lack of computing budget [^1] As we carefully fine-tune every baseline and run every experiment 15 times, this demands a significant time investment. we leave the investigations of other tasks such as those considered in Chen et al. (2021) to future work.
4 Data augmentation methods
While DA methods are continuously being proposed, objective evaluation is difficult due to the lack of extensive comparison studies. We point the reader to Chen et al. (2021) for a literature review on DA methods, and Ding et al. (2024) for one specific to generative approaches. Here, we compare LLM based methods to previous (classical) approaches. All strategies are illustrated in Figure 1. Code and all hyperparameters are available in the additional material.
4.1 Classical methods
Kumar et al. (2021); Piedboeuf and Langlais (2023) compared several “classical” DA methods, from which we select three families of strategies that have been shown to be efficient. Concretely, we test word-manipulation methods (EDA, AEDA), conditional contextual based methods (CBERT, CBART), and paraphrase based methods (T5, BT).
EDA (multiple operations on words) and AEDA (insertion of punctuation) can be implemented simply and are resource efficient. Following experiments and results from the literature, we affect 10% of the words of a sentence in EDA, and use the formula given in (Karimi et al., 2021) to calculate the number of punctuation signs to insert for AEDA. [^2] EDA works by randomly selecting one of four operations (insertion of related words, word swapping, word deletion, and word substitution) and applying it to a percentage of the words of the sentence. AEDA works by simply inserting random punctuation (among ”?”, ”.”, ”;”, ”:”, ”!”, and ”,”) into the sentence.
CBERT and CBART are more involved methods that leverage the masked words prediction task to generate new words conditionally on the class. Concretely, we mask words, prepend the class to the sentence (with a separation token), and fine-tune the model to predict the masked words. Generation of sentences follow the same process. The difference between CBERT and CBART is that latter can predict spans of words, allowing it more flexibility in the generated sentences.
BT and T5-Tapaco aim to produce paraphrases of the original sentences to bring diversity to the training set. In BT, paraphrases are generated by translating the sentence into a second language and then back into English, and we use WMT [^3] https://huggingface.co/facebook/wmt19-De-en with German as a pivot language, which has shown good performances in the past Edunov et al. (2018). T5-Tapaco makes use of T5-small-Tapaco [^4] https://huggingface.co/hetpandya/t5-small-tapaco, which is a T5 model fine-tuned on the TaPaCo paraphrase corpus (Scherrer, 2020), allowing us to directly generate paraphrases of the sentences.
4.2 Large Language Models
LLMs are ubiquitous in NLP, and have rapidly gained traction in data augmentation. Here, we test three standard strategies : paraphrasing sentences (P), zero-shot generation (ZS) by giving the LLMs a description of the task, and a 3-shot (3S) generation strategy where we give examples of the given class in addition to the description. Exact prompts used are provided in appendix B. We test two models, ChatGPT and Llama2 (using LLama-2-13B-Chat) with the same prompts.
Note that we only use LLMs to generate new data with which we train a classifier, as with any other method we compare. Using LLMs directly to label examples may lead to better classification for some tasks, but we leave this for future investigations. In any case, training a classifier is a practical solution for many problems of interest, as well as a cheaper option to deploy.
Figure 1: Strategies tested in this paper. Green algorithms are Contextual-based methods, yellow are paraphrasing methods, red are word-manipulation methods, and purple are methods using LLMs.
4.3 Baselines
Finally, we implement three simple baselines. The first one consists in training the network without the use of DA (denoted “Baseline”). The second one is an idealized strategy (denoted “‘Perfect”) where we fetch additional unused data from the training set to act as generated data. This gives an idea of the results obtainable should one collect and annotate data instead of using DA.
The last baseline is a strategy which we denote “Copy”, and which artificially inflates the size of the training set by copying multiple times (according to the ratio parameter) the original data, without modification. If the only effect of DA is — as we suspect — to help fine-tune the network better, and the modifications brought by DA are not helpful for that, then the Copy strategy should be just as efficient as all other classical DA methods.
5 Experimental setups
In this section we describe the experimental setups for our two main experiments, that is to say the testing of DA methods with better fine-tuning, and the development of better experimental protocols for DA. Our focus in this paper is data augmentation for small data, for which we use starting size of 10 and 20, but we also report results for medium training sizes (500 and 1000) in Appendix C, a setting commonly explored in the literature. When subsampling to create the training set, we make sure to choose an equal number of data points for each class, by sampling more data if needed (e.g. the actual dataset size for TREC6 and starting size of 10 is 12; two sentences per class). There is no consensus regarding the ratio of generated-to-genuine examples, but we found in both the literature and experiments that smaller datasets benefited more from larger ratios, while larger for larger datasets, we did not observe improvements with ratios larger than one. Thus, we use a ratio of 10 for the small data settings and of one for the larger dataset sizes. [^5] Based on our main results, we can hypothesize that this is simply due to pretrained models needing more time to be fine-tuned correctly on smaller datasets than medium or large ones, and the larger ratio leads to a larger number of batches, meaning more training time.
5.1 Better fine-tuning
To define our training protocol, we look at the literature to understand what is usually done. Piedboeuf and Langlais (2023) use variable patience [^6] The number of epochs during which the validation performance doesn’t increase before we stop the training process, and we then select the best model from previous epochs. for fine-tuning, depending on the dataset and dataset size, between 5 and 20 epochs, and Kumar et al. (2021) use 100 epochs of warmup followed by 8 epochs of fine-tuning from which they select the best model. Wei and Zou (2019) use early stopping with patience of 3 epochs on CNN and RNNs.
In contrast, we use a patience parameter of 50, using the validation set to find when to stop, and we fine-tune the learning rate and use label smoothing Szegedy et al. (2015) (a mean of introducing noise to the labels for regularization) before applying data augmentation, using grid search. [^7] While label smoothing was introduced in this paper as a mean of helping better learning of multiclass tasks, preliminary ablations experiments point that it raises the performance of both augmented and non-augmented dataset equally, and that the conclusions of this paper would be the same with only changing the number of epochs and hyper-parameter tuning. We report accuracy for binary tasks and the macro-F1 for multiclass tasks.
5.2 More realistic uses of data
In academic papers, DA is often evaluated with the assumption that we have access to clean validation data, often in larger quantity than what is available for the training set. This validation data is used to fine-tune both the DA algorithm and the classifier, resulting in unrealistic settings for practitioners, who may not have access to validation data or want to use their data in better ways.
We aim to test more realistic settings with the validation data, inspired by Kim et al. (2022); Zhu et al. (2023), who study the same problems but in different settings (semi-supervised learning and weakly supervised learning) and then discuss the significance of the results for textual DA. While Kim et al. (2022); Zhu et al. (2023) findings are relevant to ours, our interest in the results differ. In both papers, the authors attempt to find the best way to use the available data. In our case, we test several settings we think are more realistic, but or interest is to know if DA useful when there is little or no validation data for fine-tuning the augmentation method and the classifier.
Figure 2: Graphic representation of the four settings we test for data augmentation on small data learning. Blue represent the original training set, purple the validation set, and yellow, the test set.
We redefine the train/val. split, using different strategies illustrated in Figure 2. Firstly, we assume the validation data is not available (noVal), and we only have training data. In this setting, we train for a random number of epochs between 50 and 150, which is the range for which our models performed best in small data learning. This is inspired by the protocol of Zhu et al. (2023), where they randomly select a set of hyperparameters when no validation data is available. Secondly, we assume the presence of validation data but redefine the train/val. split, by either 1- using it all as training data and training for a random number of epochs between 4 and 8 (valAsTrain), or 2- keeping some of it as validation data (10, 250) to fine-tune both DA algorithms and the classifier (10 val/250 val).
6 Experiments
In what follows, we add the augmented material to the (small) training set and fine-tune a BERT-Base model [^8] We leave the study of other classifiers for future work. in a supervised way for each classification task, a solution which has been demonstrated efficient Devlin et al. (2019).
6.1 On the need to better fine-tuning
We first show the inefficiency of the fine-tuning protocol used in past DA studies by reporting the results they obtain without DA and comparing it to our results (without DA as well). We attempt to replicate the starting sizes, changing only the fine-tuning protocol to see the difference a longer patience, label smoothing, and better fine-tuning does. Table 2 presents the results of our experiments compared to those from the literature. Since the use of label smoothing depends on the dataset, we integrate its use as a hyperparameter.
Reported Ours
Kumar+Ubani SST-2 52.9 60.6
Kumar+Ubani SNIPS 48.6 66.8
Kumar+Ubani TREC6 79.4 91.6
Piedboeuf SST-2 87.7 87.7
Piedboeuf Irony 65.6 67.0
Piedboeuf IronyB 42.4 43.9
Piedboeuf TREC6 81.0 84.5
Table 2: Fine-tuning comparison from the literature and in this paper. All results are from training BERT without data augmentation, showing the difference made by using a longer patience as well as label smoothing.
SST2 Irony IronyB Trec6 SNIPS Average
Baseline 62.4/67.5 54.1/61.9 25.8/29.8 34.4/43.3 78.5/80.4 51.1/56.6
Perfect 82.3/85.8 62.4/58.7 19.3/33.7 72.5/80.6 94.0/94.9 66.1/70.7
Copy 61.0/68.1 54.4/57.1 25.5/30.3 36.4/49.3 77.7/83.4 51.0/57.6
EDA 63.4/69.0 54.9/58.9 23.8/30.1 35.8/48.0 76.6/83.2 50.9/57.8
AEDA 62.4/67.1 55.4/56.9 23.4/27.4 33.2/45.7 78.0/84.0 50.5/56.2
BT 60.2/66.8 55.1/58.0 24.7/27.9 29.1/51.5 75.0/82.6 48.8/57.4
CBERT 62.5/65.4 56.3/58.5 23.2/28.3 36.4/44.5 76.3/83.4 50.9/56.0
CBART 62.5/65.6 54.5/57.7 23.6/29.0 36.9/47.2 77.0/84.5 50.9/56.8
T5 64.1/67.4 56.3/58.3 24.1/28.1 34.3/48.5 76.7/84.1 51.1/57.2
GPT3.5-P 65.0/68.3 56.6/57.5 22.7/28.3 34.0/40.5 80.2/84.7 51.7/55.9
GPT3.5-ZS 82.0/78.7 56.4/56.1 23.5/24.6 40.0/48.1 84.8/88.7 57.3/59.2
GPT3.5-3S 87.7/76.1 49.2/53.6 20.9/24.6 58.4/63.9 87.2/89.6 60.7/61.5
Llama2-P 66.3/68.6 55.3/57.1 24.6/24.6 33.4/41.6 78.1/82.8 51.6/55.0
Llama2-ZS 74.8/74.0 54.3/57.2 22.2/24.1 37.3/47.3 78.1/84.1 53.4/57.3
Llama2-3S 63.3/64.4 54.1/54.8 21.0/27.2 48.6/55.8 83.3/88.2 54.0/58.1
Table 3: Average metric over 15 runs for the training set sizes of 10 (left) and 20 (right) with a ratio of 1. We report accuracy for binary tasks and macro-f1 for multiclass ones. STDs are between 0.6 and 3.0, depending on the dataset. Results for which the difference with the baseline was found to be statistically significant according to a t-test are underlined.
As we can see, carefully fine-tuning the classifier increases its performance significantly, gaining between 2 and 10%. Those gains are actually larger than the ones reported by authors while deploying DA. This aligns well with our hypothesis that the gains observed in previous studies stem only from giving more time for the model to learn. Further, if the DA strategies involve simple transformations of the sentences, it is likely that BERT is not learning new information from these transformations, and that the same result will be achieved with the Copy strategy, which we analyze in the next section.
6.2 On the inefficiency of classical DA
We now turn to DA algorithms, showing that better fine-tuning leads to essentially useless (classical) DA algorithms. Indeed we argue that alleged gains from DA that were shown in the literature can be explained by inadequate training of the (initial) classifier. Results are presented in Table 3.
For the small data setting, we notice that, on average, only two strategies really outperform the baseline : zero-shot and 3-shot generation (both using LLMs), 3-shot generation outperforming zero-shot, but not on all datasets. We explore this phenomenon in Section 6.4. The results for medium sizes (500 and 1000) are presented in Appendix C, and are concordant with more recent literature showing a lack of improvement on such a setting.
An important thing to note is that while some algorithms perform statistically better than the baseline at times, so does the “Copy” strategy. As the duplication of data should bring no new information, this suggests that the training protocol in previous studies was suboptimal and that the gains observed are not due to data augmentation.
One avenue that could be explored to bridge the gap would be to use sampling with replacement while fine-tuning BERT, which should be equivalent to duplication of data. [^9] There would still be minor differences as sampling with replacement would not represent each sample equally, but this difference should reasonably not have a large impact.
Figure 3 provides the results of the statistical tests we conducted. To simplify reading, we group together all tests between the same algorithms, meaning all tests between algorithm A and algorithm B are compiled together. Each entry in the heatmap is therefore composed of 10 t-tests (5 datasets * 2 starting sizes), and we report the percentage of t-tests for which distributions were found to be statistically different, counting only the entries for which the row algorithm performs better than the column algorithm. This gives, for each row, the measure of how many times it beats the algorithm in the column by a statistically significant margin. [^10] We provide the code along with the full saved results for all experiments, as well as the code necessary to run all statistical tests. See Figure 4 in Appendix for medium data settings (500/1000).
Baseline Perfect Copy EDA AEDA BT CBERT CBART T5 GPT3.5-P GPT3.5-ZS GPT3.5-3S Llama2-P Llama2-ZS Llama2-3S
Baseline 0 20 10 10 10 30 10 10 10 20 20 40 10 30 30
Perfect 70 0 70 70 70 70 70 70 70 70 70 80 80 80 80
Copy 20 10 0 0 10 20 10 0 0 30 10 40 20 20 30
EDA 20 10 20 0 10 20 0 0 0 10 20 40 20 10 20
AEDA 10 10 0 0 0 20 0 0 0 10 0 30 20 0 0
BT 20 10 0 0 10 0 10 10 0 10 10 30 10 10 20
CBERT 20 10 10 0 10 10 0 10 0 10 0 20 10 0 20
CBART 10 10 0 0 0 10 0 0 0 10 0 30 10 10 10
T5 20 10 0 0 0 20 10 0 0 10 0 40 10 10 20
GPT3.5-P 30 10 10 20 10 40 20 20 10 0 10 30 10 20 20
GPT3.5-ZS 70 10 60 50 50 50 60 50 50 60 0 40 60 50 30
GPT3.5-3S 60 10 60 60 60 60 60 60 60 60 40 0 60 60 50
Llama2-P 20 10 10 0 10 20 20 10 0 0 0 30 0 0 10
Llama2-ZS 30 10 20 20 20 30 20 20 20 30 0 20 40 0 30
Llama2-3S 40 0 40 40 40 40 40 40 40 40 20 20 40 40 0
Figure 3: Percent of times the row algorithm performs statistically better than the column algorithm, with a p-value threshold of 0.05 and using a two-tails paired t-test, and across the two small data settings (10/20).
From the statistical tests, we note that only three methods outperform the others by a significant and consistent margin : the perfect strategy, in which we simulate collecting and annotating external data, and the zero/k-shot generation with LLMs. All others perform equivalently and do not beat the baseline most of the time, or at least not by a significant margin. This suggests that the slight random variations we observe, which could be taken as a sign that some algorithms perform better than others, are due to randomness of the network and not a superior performance.
6.3 On more realistic uses of data
As mentioned, the second point of interest is to evaluate DA in a more realistic setting with regard to the use of validation data. We select the following algorithms : Copy, Perfect, EDA, AEDA, BT, CBERT, BART, T5, GPT3.5-ZS, and GPT3.5-3S. We use this limited selection for efficiency reasons, but also because our goal here is to establish whether DA helps at all, rather than knowing which DA technique works best. Finally, to keep in line with other DA studies, we use a ratio of 10 when the final training size is small (noVal) and of 1 for the other settings. Results are presented in table 4, and several important observations can be made.
noVal valAsTrain 10 val 250 val
Baseline 47.9 75.8 64.6 64.7
Perfect 68.8 82.4 69.4 69.9
Copy 49.6 78.6 66.3 65.8
EDA 49.1 77.8 66.0 64.9
AEDA 49.4 77.6 66.0 66.3
BT 44.8 75.9 64.6 65.0
CBERT 49.7 76.8 65.3 63.9
BART 50.6 76.6 65.3 63.9
T5Par 49.4 77.2 65.3 64.6
GPT3.5-ZS 55.9 76.9 65.6 64.4
GPT3.5-3S 57.2 71.6 67.2 64.8
Table 4: Average metrics over the five datasets for different settings and DA strategies. The best result for each setting is in bold.
First, in almost all cases and contrary to the main results of this paper, data augmentation reveals itself to be useful. However, the fact that the “Copy” strategy often outperforms the other strategies reinforces our belief that the use of data augmentation is simply to facilitate fine-tuning the network.
Second, there is no difference in performance between having 10 and 250 data points as validation data, with the rest as training set. This could be related to diminished increase in model performance as dataset size augments, or point to a trade-off in validation/training size and that a validation set of 10 data points is too small to be of use.
Third, we confirm here that methods adding something akin to external data (GPT3.5-ZS and GPT3.5-3S) are the most useful techniques, but only on small training sizes. On larger ones, those methods do not bring any improvement whatsoever, most likely due to the general inefficiency of data augmentation in this setting. [^11] This is concordant with the literature and our own experiments. We also note that while the performance of GPT3.5-ZS is higher than GPT3.5-3S for two settings, this is due exclusively to its poor performance on the Irony datasets, which we discuss in Section 6.4. For the other datasets, GPT3.5-3S outperforms GPT3.5-ZS by a large margin.
Finally, it seems that the best strategy overall when the total amount of available data is only a few hundred is to use all of them as training data, with the Copy strategy. [^12] Or, we can assume, a larger number of epoch. This outperforms by a large margin the strategies where we keep some data as validation data, most likely because these strategies overfit. As mentioned, however, there is a large gap between 10 and 250 data and it is possible that some amount of data in the middle would perform better. We leave this investigation for future work.
6.4 Analysis of DA with LLMs
Our main results show that in a small data setting, 3-shot generation was often more efficient than zero-shot generation, by as much as 18% in some cases. However, this does not hold for the two irony datasets, for which zero-shot outperforms 3-shot. In this section, we analyze the output of the algorithms, in the hope of understanding better why some prompt mechanisms work well on some datasets but not on others.
We first show examples of generated sentences in Table 5 and in Appendix A. For the sake of space, we only show generated sentences for the tasks of SST-2 and Irony, but we give all sentences generated by ChatGPT in the associated Github.
Couldn’t have asked for a better night.
-P Could not have hoped for a more perfect evening.
Non-Ironic -ZS Wow, I absolutely love it when my phone battery dies in the middle of an important call. Just what I needed!
-3S Enjoying a relaxing day at the beach with clear skies and warm sunshine. #PerfectDay
Work should be fun today :unamused_face:
-P Today at work is expected to be unenjoyable :unamused_face:
Ironic -ZS Oh, how wonderful! Another day of unexpected surprises and setbacks. It’s like the universe’s way of keeping me on my toes!
-3S Declared I’d be more organized and lost my keys for the umpteenth time. I’m the queen of tidiness! :key::mag: #OrganizationQueen
Table 5: Generated sentences by the GPT3.5-based DA methods for the Irony dataset for one ironic and one non-ironic example.
By observing generated sentences, we see that ChatGPT has difficulty understanding the concept of Irony, as well as sticking to the more familiar language of Twitter. As such, adding more sentences doesn’t help because the new sentences are too far from the training distribution to bring valuable information. Nevertheless, the 3-shot strategy does seem to bring the generated sentences closer to the training distribution, as we can at least observe that the new sentences contain hashtags and emojis. It is likely that by further fine-tuning the prompts, we would reduce the gap between the data distribution and the generated sentences and generate highly informative sentences.
7 Conclusion
In this paper, we test DA for sentence classification and show that in both medium and small data learning, performances of DA had been overestimated by the use of inadequate training protocols. Furthermore, while we have shown that classical DA methods are inefficient, we have also demonstrated that simulating external data collection with LLMs does improve the performance, and future work should therefore focus on this rather than techniques modifying genuine sentences.
We want to emphasize that our results are only looking at balanced classification tasks, and furthermore only at short text classification using BERT as a classifier. While this seems a small field of study, it is one of the most popular in textual DA literature, making our findings significant.
Future work should also evaluate whether the protocols are adapted for fields close to interpretable DA. Notably non-interpretable DA, or DA for unbalanced data, might still be an efficient mean of increasing performances. Other textual tasks might also still benefit from DA, due to different goals. As an example, data augmentation for the task of keyphrase generation (Ray Chowdhury et al., 2022; Garg et al., 2023) aims to encourage the network to generate more keyphrases that are absent from the input. It is very plausible that data augmentation may be useful here.
Our findings are important as they answer many questions in textual DA which have been standing for years, namely which data augmentation algorithms is best (they all perform similarly except for data generation), why does it help (gives more time to the network to learn), and what makes a generated sentence informative (they don’t bring contradictory information to the network). As noted, data generation remains one of the best way to perform data augmentation, if one can get the LLMs to generate data adequate for the training distribution.
8 Limitations
As already stated, this study has a specific, although important, scope. We didn’t look at the impact of data augmentation on other languages than English, or other textual tasks, and while we performed fine-tuning to the best of our knowledge, it is possible that some combinations of hyperparameters that were not explored may change some of our conclusions. We also omitted an ablation study of the fine-tuning (patience, label smoothing, grid search) due to time limitation, but this would provide more insight into the mechanism of training BERT. Furthermore, our exploration of hyperparameters for ChatGPT (mainly the temperature) was limited, and we only used the Llama2-13b-chat model as a comparison, but a comparison of other sizes and also other LLMs should be included in future work, such as Llama3 Meta (2024), Gemini Team et al. (2024), GPT-4 OpenAI et al. (2024), and others. Finally, we only looked at BERT as a classifier, which is the most common one used in DA studies, but it would be pertinent to not only look at other pretrained transformers but also at the impact of DA on more classical models such as CNNs, RNNs, or even statistical methods (SVM, NB, etc).
On a broader scope, there is no guarantee our results hold for other textual tasks (such as question answering, explainability, or keyphrase generation), which have different structures. A final point is that we only compared ourselves to the studies of Kumar et al. (2021); Piedboeuf and Langlais (2023) to demonstrate the inefficiency of fine-tuning in past studies. While these are some of the most complete we found, it would be pertinent to repeat the experience with other papers.
Appendix A Generated sentences
Negative Dataset apparently reassembled from the cutting-room floor of any given daytime soap.
GPT3.5-P It appears to have been pieced together from discarded scenes of any typical daytime soap opera.
GPT3.5-ZS The film was a major disappointment, lacking any coherent plot or engaging characters.
GPT3.5-3S It’s a shame that such a promising concept was wasted on a poorly executed film.
Positive Dataset béart and berling are both superb , while huppert … is magnificent.
GPT3.5 - P Both Béart and Berling deliver exceptional performances, while Huppert… she is truly magnificent.
GPT3.5-ZS This film is a refreshing breath of fresh air, offering a unique and original take on familiar themes.
GPT3.5-3S This film is a delightful and heartwarming celebration of friendship, full of laughter and touching moments.
Table 6: Examples of generated sentences for the SST-2 dataset.
Appendix B Prompts
In this section, we describe the prompts we used for ChatGPT and LLama2. For the paraphrasing strategy, we copy Piedboeuf and Langlais (2023), who simply asks for paraphrases in batch if the ratio is 1 and for multiple paraphrases if the ratio is larger than one. We refer to their paper for more detail on the prompting.
For the zero and three-shot generation, we use the template shown in Table 7.
DATASET_DESC CLASS_DESC
movie reviews negative or somewhat negative
positive or somewhat positive
headline Fake/Real news classification Real
Fake
Ironic tweet detection Non-Ironic Tweets
Ironic Tweets
Ironic tweet detection Tweets that are not ironic
Tweets ironic by polarity contrast, where the polarity is inverted between the literal and intended evaluation
Tweets ironic by Situational Irony, where a situation fails to meet some expectation
Tweets ironic by Other type of Irony, where the Irony is neither by Polarity Contrast or by Situational Irony
Question classification Questions about an abbreviation
Questions about an entity (event, animal, language, etc)
Question concerning a description (of something, a definition, a reason, etc)
Questions about a human (description of someone, an individual, etc)
Questions about a location
Questions about something numerical (weight, price, any other number)
Table 7: Prompt patterns for the zero-shot strategies for ChatGPT and Llama. The prompt is of the form “Here are some examples of \$CLASS_DESC from a dataset of \$DATASET_DESC: \$EXAMPLES. Can you create 10 more sentences of that class that would fit the dataset ”.
We referred to the description given in the original papers of each dataset to craft informative prompts.
Appendix C Supplementary Results
SST2 Irony IronyB TREC6 SNIPS Average
Baseline 87.7/88.8 67.0/68.1 43.9/45.9 84.5/87.9 95.6/96.2 75.8/77.4
Perfect 88.9*/89.7* 67.3/71.1* 46.0*/48.5* 87.6*/89.3 96.1*/96.9* 77.2/79.1
Copy 87.7/88.8 66.2/69.2 43.7/46.7 83.8/86.3 95.5/96.4 75.4/77.5
EDA 88.1/88.9 65.9/69.4 43.6/47.2* 82.5/86.1 95.8/96.3 75.2/77.6
AEDA 88.0/89.0 65.1/69.5 42.8/46.7 82.7/87.4 96.1*/96.3 74.9/77.8
BT 88.0/89.2* 64.9/68.4 44.2/45.8 83.1/87.1 95.5/96.3 75.1/77.3
CBERT 87.6/88.6 64.8/69.5 43.6/46.1 82.9/84.7 95.4/96.3 74.9/77.0
CBART 87.7/88.7 66.0/68.8 43.6/45.1 81.2/82.5 95.9/96.2 74.9/76.3
T5 87.5/88.8 64.9/67.7 43.6/45.6 82.9/84.8 95.8/96.3 75.0/76.6
GPT3.5-P 88.2*/89.3* 65.1/69.5* 44.3/47.5 84.5/86.8 95.5/96.1 75.5/77.8
GPT3.5-ZS 87.7/88.8 64.2/67.4 42.0/45.1 81.9/86.0 95.5/95.8 74.3/76.6
GPT3.5-3S 88.1*/88.6 64.9/69.0 40.4/44.4 83.0/86.8 95.9/96.2 74.5/77.0
Llama2-P 87.8/88.9 65.4/68.1 44.1/46.1 83.3/85.9 95.7/96.4 75.3/77.1
Llama2-ZS 88.1*/88.8 65.0/69.2 42.5/45.7 82.2/86.0 95.6/96.0 74.7/77.2
Llama2-3S 87.8/88.4 65.0/67.9 42.8/46.2 83.8/85.3 95.7/96.4* 75.0/76.8
Table 9: Average metric over 15 runs for the training set sizes of 500 (left) and 1000 (right) with a ratio of 1. We report accuracy for binary tasks and macro-f1 for multiclass ones. STDs are between 0.6 and 3.0, depending on the dataset. Stars represent results for which the difference with the baseline was found to be statistically significant.
Baseline Perfect Copy EDA AEDA BT CBERT CBART T5 GPT3.5-P GPT3.5-ZS GPT3.5-3S Llama2-P Llama2-ZS Llama2-3S
Baseline 0 0 0 20 10 10 30 20 20 0 40 30 20 30 30
Perfect 80 0 90 70 70 90 100 80 90 80 100 90 90 100 90
Copy 0 0 0 0 0 0 0 30 10 0 40 20 0 10 20
EDA 10 0 0 0 0 0 0 20 20 0 50 20 10 10 10
AEDA 10 0 0 10 0 10 20 20 20 20 40 30 20 20 30
BT 10 0 0 0 0 0 20 20 10 0 20 20 0 20 10
CBERT 0 0 0 0 0 0 0 10 10 0 30 20 10 0 10
CBART 0 0 0 0 0 0 10 0 0 0 20 10 0 0 0
T5 0 0 0 0 0 0 0 10 0 0 20 10 0 10 0
GPT3.5-P 30 0 20 20 10 10 30 40 50 0 60 30 0 40 20
GPT3.5-ZS 0 0 0 0 0 0 0 10 0 0 0 10 0 0 0
GPT3.5-3S 10 0 10 0 0 0 10 10 10 10 30 0 0 0 0
Llama2-P 0 0 0 0 0 0 10 10 0 0 30 20 0 10 10
Llama2-ZS 10 0 10 0 0 0 0 10 10 0 20 10 10 0 20
Llama2-3S 10 0 0 0 0 0 10 10 0 10 20 20 0 10 0
Figure 4: Percentage of times the row algorithm performs statistically better than the column algorithm, with a p-value threshold of 0.05 and using a two-tails paired t-test and with the starting sizes of 500/1000.